OpenStudy (anonymous):

A function f is defined by:....(I have to use the equation thing to type it out)

OpenStudy (anonymous):

\[f:x \rightarrow \frac{ 1 }{ x } \]

OpenStudy (anonymous):

\[x \in \mathbb{R}, x \neq 0\]

OpenStudy (anonymous):

Find \[f ^{2}(x)\]

OpenStudy (anonymous):

I am not sure how to do this, please help!

OpenStudy (amistre64):

replace f(x) by its definition ...

OpenStudy (amistre64):

since f(x) = 1/x what is the square of f(x)?

OpenStudy (anonymous):

\[\frac{ 1 }{ x^2 }\] ?

OpenStudy (jango_in_dtown):

right